OpenStudy (anonymous):

Find the area of an equilateral triangle with sides of length 5.

OpenStudy (amistre64):

5^2 sin(60)/2

OpenStudy (amistre64):

sin(60) = sqrt(3)/2 if you just wanna sub it in there

OpenStudy (anonymous):

if you don't know trignometry..then draw perpendicular and use pythogeros theorem to find length of the perpendicular and use half base cross height!

OpenStudy (anonymous):

if you have learn't trig.. then don't bother seeing my replies :D

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

The area of an equilateral triangle where each side is equal to "a" in length is:\[\frac{\sqrt{3} a^2}{4} \]When a is equal to 5 the area is:\[\frac{5^2\sqrt{3} }{4}=10.8253 \]